OpenStudy (anonymous):

Use these points to write the equation of the parabola in standard form (3,0) (-4,-5) (4,-5)

ganeshie8 (ganeshie8):

say the required parabola eq'n is : \(a(x-h)^2 + k\)

ganeshie8 (ganeshie8):

it goes thru two symmetrical points about y-axis : (-4,-5) and (4,-5) that means, x-coordinate of its vertex is 0 h = 0

OpenStudy (anonymous):

Yeah that's what I got I was trying to make sure thank you

ganeshie8 (ganeshie8):

find other values a and k, by plugging in any two points

ganeshie8 (ganeshie8):

\(a(x-0)^2 + k\) plugin (3, 0) above 0 = a(3)^2 + k 0 = 9a + k --------(1) plugin (4,-5) above -5 = a(4)^2+k -5 = 16a + k --------(2) solving both, a = -5/7 k = 45/7

ganeshie8 (ganeshie8):

so the required parabola wud be :- \(\frac{-5}{7}(x-0)^2 + \frac{45}{7}\)

OpenStudy (anonymous):

Yay! I was right thank you

ganeshie8 (ganeshie8):

np... u wlc :)

OpenStudy (anonymous):

Another question, sorry for being a bother for Vertex form I use the equation of f(x)=ax^2+bx+c right?

ganeshie8 (ganeshie8):

thats standard form !

ganeshie8 (ganeshie8):

standard form : f(x) = ax^2+bx+c vertex form : f(x) = a(x-h)^2+k